OpenStudy (anonymous):

Any of you familiar with Avogadro Goes to Court- Case Study? I have to calculate the cost in dollars of a single aluminum atom in the roll of aluminum foil. The roll cost is $1.59, and the area of the roll is 6.96 square metres. And I am given approximately 20.0 cm by 20.0 cm (8 inch by 8 inch) section of the roll that I have cut out.

OpenStudy (anonymous):

what is mass of that square 20x20 you cut out?

OpenStudy (anonymous):

2.5838g

OpenStudy (anonymous):

i got that price of one atom is 1,58*10^-25 $

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

ok lets start: you are given that 6,96 m2 has price of 1,59 $ now you have cut out 20x20 piece which has surface area of 400 cm2 and 400 cm2 = 0,04 m2 now share is: 0,04/6,96= 5,747*10^-3 and share of price is: 5,747*10^-3 * 1,59$ = 9,1379*10^-3 $ now you take mass of that 20x20 piece and calculate amount: n=m/M n= 2,5838g/ 26,9815 gmol-1=0,095762 mol now lets calculate number of atoms that corresponds: n=N/L --> N= n*L N= 0,095762mol * 6,022*10^23 mol-1 = 5,76678*10^22 now you must calculate price of 1 atom so lets again calculate share: 1/5,76678*10^22 = 1,734*10^-23 and now lets calculate what share of price that corresponds: 1,734*10^-23 * 9,1379*10^-3 = 1,58*10^-25 $

OpenStudy (anonymous):

how do i find the total amount of the aluminum for 8inch by 8inch then?

OpenStudy (anonymous):

it is calculated there are 5,76678*10^22 atoms in 20x20 piece

OpenStudy (anonymous):

oh thank you very much!!!!!

OpenStudy (anonymous):

not a problem, glad to help